• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 00:18:45 on December19,2003

    Entry number 114592

    keyword=Swing shift summary

    Started shift with beam energy width measurements underway. MCC is still working on the analysis, but preliminary results indicate that at 4 GeV, the energy spread into Hall A was ~2.5E-5. Hall C beam for G0 was not achieved simultaneously at full current.

    Restored 2 GeV beam to Hall A by 9:30 for start of commissioning runplan.

    Started with 50 A in right septum, looking for beam steering. On the beam dump viewer, 50 A appears to shift the beam spot substantially. Following discussions with MCC and a test of how much the beam is shifted at target by a steering coil, we estimate the shift to be at the order of cm at 50 A.

    Following discussion with RC, MCC, RADCON and shift, we decided to keep beam current at 2 uA and ramp septum current to 100 A. Dump ion chambers tripped at about 88 A.